OpenStudy (anonymous):

Solve the following system by using graph paper or graphing technology. 3x – 2y = 7 9x – 6y = 3 What is the solution to the system? No Solutions Infinite Solutions (3, 2) (3, –2)

OpenStudy (anonymous):

manipulate the second equation : y = 3/2x -1/2 then substitute into first equation, or times the first equation by -3 to get -9x + 6y = -21 to eliminate (x) and solve for y.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no solutions

OpenStudy (anonymous):

@strenesmee you shouldn't just say an answer

OpenStudy (anonymous):

ok -9x+6y=-21 9x-6y=3 ox=-18, so there is no solutions is it better now??

OpenStudy (anonymous):

The graph of a system of equations with the same slope and the same y-intercepts will have no solutions. Always Sometimes Never

OpenStudy (anonymous):

joshdavoll, whats wrong with him just saying an answer?